Calculate the molar solubility of Mg(OH)2 in the following solvents. Ksp = 1.8 x 10¯11

pure water

8.68×10?2 M MgCl2

3.65×10?2 M KOH(aq).

Really i just need someone to explain how these things affect solubility numerically
Answer #1

Mg(OH)2\rightleftharpoons Mg2+ + 2OH-

In pure water,

Ksp = [ Mg2+] * [OH-]2   

= x*(2x)2 = 4x3 =  1.8 x 10-11 => x = 1.65*10-4mol/lt

In MgCl2,

Mg(OH)2 Mg2+ OH-
Initial 8.68*10-2 ( Already MgCl2 has Mg2+ ions) 0
Change +x +2x
Equilibrium 8.68*10-2 + x 2x

Ksp = [ Mg2+] * [OH-]2   

= (8.68*10-2 + x)(2x)2 = 1.8*10-11 as ksp is very low (8.68*10-2 + x) = 8.68*10-2

=> 8.68*10-2*4x2 = 1.8*10-11

Solving, x = 7.2*10-6M

In KOH,

Mg(OH)2 Mg2+ OH-
Initial 0 3.65*10-2 (( Already KOH has OH- ions)
Change +x 3.65*10-2+2x
Equilibrium x 3.65*10-2+2x

Ksp = [ Mg2+] * [OH-]2   

= ( x)(3.65*10-2+2x)2 = 1.8*10-11 as ksp is very low (3.65*10-2+2x) = 3.65*10-2

=> x*(3.65*10-2)2 = 1.8*10-11

Solving, x = 1.35*10-8M